### v3

# On the connectedness of the singular set of holomorphic foliations

Abstract.
Let $\mathscr{F}$ be a singular holomorphic foliation of dimension $k>1$ in a projective $n$-manifold $X$. Assume that the normal sheaf of $\mathscr{F}$ is ample (as is always the case when $X=\mathbb{P}^{n}$) and the singular set $\mbox{Sing}(\mathscr{F})$ has dimension $\leq k-1$. We show that the union of the irreducible components in $\mbox{Sing}(\mathscr{F})$ of dimension exactly k1k-1 is necessarily connected. Consequently, this yields a Bott-type topological obstruction to the integrability of singular holomorphic distributions, echoing Bott’s vanishing theorem, and it answers a question by D. Cerveau for codimension one foliations on $\mathbb{P}^{3}$.
Introduction
In the 1960s, Raoul Bott [3] demonstrated that topological constraints can obstruct the existence of globally integrable foliated structures. In the same vein, we establish a global obstruction to integrability for singular holomorphic distributions. Indeed, there exist holomorphic distributions of dimension $k>1$ on $\mathbb{P}^{n}$, for all $n\geq 3$, whose singular set contains several connected components of dimension k1k-1 (see Remark 4.5). Our main Theorem precludes integrability for all such distributions.
We recall [2] that a $k$-dimensional singular holomorphic foliation on a complex manifold $X$ of dimension $n$ is defined by an exact sequence
$0\to\mathscr{F}\longrightarrow TX\longrightarrow N_{\mathscr{F}}\longrightarrow 0\,$,
where $TX$ is the sheaf of holomorphic vector fields on $X$, the tangent sheaf $\mathscr{F}$ is an integrable coherent subsheaf of $TX$ of rank $k$, and the normal sheaf $N_{\mathscr{F}}$ is coherent and torsion free, of rank nk;n-k; this is the codimension of $\mathscr{F}$. The singular scheme, denoted by ${\rm Sing}(\mathscr{F})$, is the locus where the normal sheaf is not locally free. This implies that ${\rm Sing}(\mathscr{F})$ has codimension $\geq 2$. Let $\det(N_{\mathscr{F}})$ denote the determinant bundle of the normal sheaf. We prove:
Theorem 1.
Let $X$ be a projective manifold of dimension $n$, and $\mathscr{F}$ a singular holomorphic foliation of dimension $k>1$ such that the bundle $\det(N_{\mathscr{F}})$ is ample. Let the singular set be ${\rm Sing}(\mathscr{F})=Z\cup Z_{-}$, where $Z$ has dimension k1k-1 and $Z_{-}$ consists of lower-dimensional components. Then $Z$ is a connected set.
Our result is sharp since all the assumptions are necessary; see Section 4.
Theorem 1 answers positively the question posed by D. Cerveau in [7] regarding dimension 2 foliations on $\mathbb{P}^{3}$, since the determinant bundle of the normal sheaf of every holomorphic foliation on projective spaces is ample. Cerveau’s problem was motivated by the known models of codimension one foliations classified in degrees 0, 1, and 2.
This falls within the Bott phenomenon of global obstructions to integrability: whereas Bott’s theorem compels the vanishing of invariant polynomials in the Chern classes of the normal bundle in degrees exceeding the codimension for integrable subbundles, our connectedness Theorem 1 furnishes a topological obstruction to integrability for singular distributions under an ampleness hypothesis. We know from [5, Theorem 9.5 item 2b.] and [13, Table 1] that there are families of holomorphic distributions of dimension two on $\mathbb{P}^{3}$, whose singular set contains two connected components of dimension 1.
Our proof uses in an essential way the Baum–Bott residues [2]. The idea is as follows. By Baum–Bott, the Chern number $c_{1}(N_{\mathscr{F}})^{n-k+1}[X]$ has a residue localized at each connected component $Z_{1},\dots,Z_{s}$ of the singular set of dimension k1k-1; these residues are elements in $H_{2k-2}(Z_{i})$. Then, the ampleness condition yields, after some work, that the residue at each $Z_{i}$ equals the total Chern number $c_{1}^{n-k+1}(N_{\mathscr{F}})\cap[X]$. Hence $s=1$.
We give examples showing that the ampleness condition on the determinant of the normal sheaf is necessary for Theorem 1. This condition is always satisfied for holomorphic distributions in $\mathbb{P}^{n}$, see for instance Remark 4.6.
We show in Section 4 that the ampleness hypothesis on $\det(N_{\mathscr{F}})$ is indispensable for Theorem 1. Moreover, when $X=\mathbb{P}^{n}$ and $n\leq 2(k-1)$, any two irreducible projective subvarieties $V_{1},V_{2}\subset\mathbb{P}^{n}$ of dimension k1k-1 must meet. Indeed, in $A^{*}(\mathbb{P}^{n})=\mathbb{Z}[h]/(h^{n+1})$ we have
$V_{1}\cap V_{2}]=\deg(V_{1})\deg(V_{2})h^{\,2n-2k+2}\neq 0.$
Consequently, the (k1)(k-1)-dimensional components of ${\rm Sing}(\mathscr{F})$ cannot be pairwise disjoint, and the conclusion of Theorem 1 is automatic in this range. However, this phenomenon does not hold in general. For example, consider $X=\mathbb{P}^{1}\times\mathbb{P}^{5}$ and subvarieties $W_{1}=\{p\}\times\mathbb{P}^{4}$ and $W_{2}=\{q\}\times\mathbb{P}^{4}$, where $p\neq q$. Here, $n=6$ and $k=4$, so n=2(k1)n=2(k-1), yet $W_{1}$ and $W_{2}$ are clearly disjoint.
In [4], the first author, Cerveau, Giraldo, and Lins Neto asked whether the tangent sheaf of a codimension one foliation must split as a sum of line bundles whenever it is locally free. Recently, the third author, Faenzi and Vallés in [12] provided a negative answer to this problem by producing examples of rational foliations on $\mathbb{P}^{3}$ of even degree $2k+4,~k\geq 0$, whose tangent sheaf is a twisted null-correlation bundle, hence non-split; the singular sets of such foliations are singular, non-reduced curves. Furthermore, these foliations admit flat deformations to non-integrable distributions.
This naturally leads to the following question:
What are the topological obstructions for a rank $2$ reflexive sheaf to be the tangent sheaf of a foliation in a projective 3-fold?
Answering the preceding question in the setting of projective threefolds, we show that a natural topological constraint on the tangent sheaf—together with the ampleness of the normal determinant—yields a global obstruction to integrability, detected by the number of isolated singularities of the foliation.
Corollary 1.
Let $\mathscr{F}$ be a codimension one foliation on a smooth projective threefold $X$ satisfying $h^{1}(\mathcal{O}_{X})=0$, and let $C$ be the union of all 1-dimensional components of ${\rm Sing}(\mathscr{F})$; assume that $C$ is reduced and that the determinant of normal bundle of $\mathscr{F}$ is ample. If $h^{1}(TX\otimes\det(N_{\mathscr{F}})^{*})=h^{2}(TX\otimes\det(N_{\mathscr{F}})^{*})=0$, then
$h^{2}(\mathscr{F}\otimes\det(N_{\mathscr{F}})^{*})=c_{3}(\mathscr{F})\cap[X]$
which corresponds to the number of isolated singularities of $\mathscr{F}$ counted with multiplicities. In particular, when $X=\mathbb{P}^{3}$, the conditions are satisfied and the formula holds.
This is also related to another problem posed by Cerveau in [7, Problem 2.2], which concerns the description of foliations whose singular set has pure dimension one, a condition equivalent to the tangent sheaf being locally free (see [5]). The result provides a topological and algebro-geometric obstruction for such foliations.
In particular, a codimension 1 foliation with locally free tangent sheaf on a smooth 3-dimensional projective manifold, under the hypotheses of Corollary 1, must satisfy
$h^{2}(\mathscr{F}\otimes\det(N_{\mathscr{F}})^{*})=0.$
This condition reflects a topological obstruction for a locally free sheaf to be the tangent sheaf of a foliation.
Moreover, codimension 1 foliations on $\mathbb{P}^{3}$ with locally free tangent sheaf force ${\rm Sing}(\mathscr{F})$ to be singular, unless $\mathscr{F}$ is a pencil of planes.
Corollary 2.
Let $\mathscr{F}$ be a codimension one foliation on $\mathbb{P}^{3}$. If the tangent sheaf of $\mathscr{F}$ is locally free, then either $\mathscr{F}$ is a pencil of planes, or ${\rm Sing}(\mathscr{F})$ is a singular curve.
We emphasize that the hypothesis that $\mathscr{F}$ is locally free is essential: recall that the singular scheme of a generic rational foliation of type $(a,b)$ consists of a smooth complete intersection curve of degree $ab$ plus
$(a+b-2)^{3}+2(a+b-2)^{2}+2(1-ab)(a+b-2)$
distinct points, see [5, Section 10.1] and [11] for the higher-dimensional case. The example in [12, Theorem 7.1], where the foliation is given by a pencil and the tangent sheaf is a twisted null-correlation, has as singular set a singular curve.
Acknowledgments
CA and JS gratefully acknowledge the University of Bari for its hospitality and support through UNIBA Visiting Professor Program. JS is partially supported by UNAM-PAPIIT grant IN101424. MC is partially supported by the Università degli Studi di Bari and by the PRIN 2022MWPMAB- “Interactions between Geometric Structures and Function Theories” and he is a member of INdAM-GNSAGA; he would like to thank Nigel Hitchin for insightful discussions during his visit to the University of Oxford. MJ is supported by the CNPQ grant number 305601/2022-9, the FAPESP-ANR project number 2021/04065-6.
1. Proof of Theorem 1
The proof uses Baum–Bott residues, and by definition, these are elements in the homology groups of ${\rm Sing}(\mathscr{F})$. Yet, it is convenient in the sequel to use these residues in cohomology, so we consider the following commutative diagram:
$H^{2(n-k+1)}(X,X\setminus Z)$$H^{2(n-k+1)}(X)$$H^{2(n-k+1)}(U_{i},U_{i}\setminus Z_{i})$$H_{c}^{2(n-k+1)}(U_{i})$$H_{2k-2}(Z)$$H_{2k-2}(X)$$H_{2k-2}(Z_{i})$$H_{2k-2}(U_{i})$$\scriptstyle{\phi}$$\scriptstyle{A}$$\scriptstyle{j^{*}_{U_{i}}}$$\scriptstyle{P}$$\scriptstyle{j^{*}_{U_{i}}}$$\scriptstyle{\phi_{i}}$$\scriptstyle{A_{i}}$$\scriptstyle{P_{i}}$$\scriptstyle{(i_{Z})_{*}}$$\scriptstyle{P^{-1}}$$\scriptstyle{(i_{Z_{i}})_{*}}$$\scriptstyle{(j_{Z_{i}})_{*}}$$\scriptstyle{(i_{Z_{i}})_{*}}$$\scriptstyle{A_{i}^{-1}}$$\scriptstyle{(j_{U_{i}})_{*}}$$\scriptstyle{P_{i}^{-1}}$
where all the cohomology and homology groups are with coefficients in $\mathbb{C}$; $P=\,\cap[X]$ is Poincaré duality; the $U_{i}$ are pairwise disjoint smooth regular neighborhoods (as in [15]) of the connected components $Z_{1},\dots,Z_{s}$ of $Z$, AA is Alexander duality, all the other homomorphisms are induced by the inclusion. The morphism $\phi$ comes from the following long exact sequence of relative cohomology
$H^{2(n-k+1)}(X,X\setminus Z)$$H^{2(n-k+1)}(X)$$H^{2(n-k+1)}(X\setminus Z).$$\scriptstyle{\phi}$
Recall that $c_{1}(N_{\mathscr{F}})=c_{1}({\rm det}(N_{\mathscr{F}}))$, see [16, Section 5.6]. The residues that we consider are localizations of $c_{1}(N_{\mathscr{F}})^{n-k+1}$ in $H_{2k-2}(Z_{i})$. Given one such residue
${{\rm Res}}_{c_{1}^{n-k+1}}(\mathscr{F},Z_{i}):={{\rm Res}}(\mathscr{F},Z_{i})\in H_{2k-2}(Z_{i}),$
we can push it (by the inclusions) to $H_{2k-2}(X)$ and lift it by the Poincaré isomorphism to $H^{2(n-k+1)}(X)$. Denote it by ${\widehat{\rm Res}}(\mathscr{F},Z_{i})$. That is:
${\widehat{\rm Res}}(\mathscr{F},Z_{i})=P^{-1}\big((i_{(Z_{i})})_{*}{\rm Res}(\mathscr{F},Z_{i})\big)\,$,
where $(i_{(Z_{i})_{*}})$} denotes the morphism $H_{*}(Z_{i})\to H_{*}(Z)\to H_{*}(X)$. Then we have the Baum–Bott formula in the cohomology group $H^{2(n-k+1)}(X)$:
$c_{1}(N_{\mathscr{F}})^{n-k+1}=\sum_{i=1}^{s}\;{\widehat{\rm Res}}(\mathscr{F},Z_{i})\,.$
First of all, we observe that the previous formula implies that the ampleness of $c_{1}(N_{\mathscr{F}})$ ensures that the component $Z=\bigcup_{i=1}^{s}Z_{i}$ of dimension k1k-1 is non-empty, since $c_{1}(N_{\mathscr{F}})^{n-k+1}\neq 0$, and component $Z_{-}$ of dimension smaller than k1k-1 does not contribute to $^{\widehat{\rm Res}}$.
Now, consider the above pairwise disjoint smooth regular neighborhoods of the connected components of $Z$. That is, for each connected component $Z_{i}\subset Z$, we take an open neighborhood $U_{i}$ with a retraction
${Z_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{Z_{i}}}$${U_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho_{i}}$${Z_{i}}$
with a smooth boundary $\partial U_{i}$, such that
$\text{for all }j\neq i\quad\overline{U_{i}}\cap\overline{U_{j}}=\emptyset,\quad(\overline{U_{i}}\setminus Z_{i})\cap Sing(\mathscr{F})=\emptyset.$
Let ${\rm Res}(\mathscr{F},Z_{i})$ be the Baum–Bott residue in $H_{2k-2}(Z_{i})$. Define:
$\widehat{\rm Res}(\mathscr{F}|_{U_{i}},Z_{i})=P_{i}^{-1}\big((i_{Z_{i}})_{*}{\rm Res}(\mathscr{F},Z_{i})\big)\,.$
By the commutativity of the diagram above, we have the following equality in $H_{c}^{2(n-k+1)}(U_{i})$:
(1.1) $\widehat{\rm Res}(\mathscr{F}|_{U_{i}},Z_{i})=j^{*}_{U_{i}}\big(\widehat{\rm Res}(\mathscr{F},Z_{i})\big)\;.$
We can take a neighborhood $V_{i}$ of $\partial U_{i}$ such that $\mathscr{F}$ is regular on $V_{i}$, i.e., the tangent sheaf of $\mathscr{F}$ is locally free on $V_{i}$, where $U_{i}$ is a retraction of $Z_{i}$. Then, we can apply Theorem 1.4 from [17] and, using the naturality of Chern classes, we obtain
(1.2) $\widehat{\mathrm{Res}}(\mathscr{F}|_{U_{i}},Z_{i})$ $=c_{1}(\det(N_{\mathscr{F}})|_{U_{i}})^{n-k+1}$
$=c_{1}(j^{*}_{U_{i}}\det(N_{\mathscr{F}}))^{n-k+1}$
$=j^{*}_{U_{i}}\big(c_{1}(N_{\mathscr{F}})^{n-k+1}\big)\;.$
Write $Z_{i}$ as the union of its irreducible components of dimension k1k-1:
$Z_{i}=\bigcup_{\alpha=1}^{m_{i}}W_{i\alpha}\qquad(\mathrm{dim}W_{i\alpha}=k-1).$
The localized residue in $U_{i}$ then reads
$\widehat{\mathrm{Res}(\mathscr{F}|_{U_{i}},Z_{i})}=\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,[\widehat{W_{i\alpha}}]\ \in H_{c}^{2(n-k+1)}(U_{i}).$
Applying $j^{*}_{U_{i}}$ we get the following equality in $H_{c}^{2(n-k+1)}(U_{i})$:
(1.3) $j^{*}_{U_{i}}\big(\widehat{\rm Res}(\mathscr{F},Z_{i})\big)\,=\,\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big([\widehat{W_{i\alpha}}]\big)\,=\,j^{*}_{U_{i}}\big(c_{1}({\rm det}(N_{\mathscr{F}})\big)^{n-k+1}\big),$
where the last equality follows from equation (1.2). The ampleness says us that here there is an embedding $i:X\hookrightarrow\mathbb{P}^{N},$ for some $N$, such that $\det(N_{\mathscr{F}})=\mathcal{O}_{X}(r)=i^{*}\bigl(\mathcal{O}_{\mathbb{P}^{N}}(r)\bigr).$ Therefore
$c_{1}({\rm det}(N_{\mathscr{F}}))^{n-k+1}=r^{n-k+1}h^{n-k+1}\in H^{2(n-k+1)}(\mathbb{P}^{n})\,$,
where $h:=c_{1}({\mathcal{O}}_{X}(1))$. Hence
$j^{*}_{U_{i}}\big(c_{1}({\rm det}(N_{\mathscr{F}}))^{n-k+1}\big)=r^{n-k+1}j^{*}_{U_{i}}(h^{n-k+1})\in H^{2(n-k+1)}(U_{i})\,.$
Since $k>1$, we can apply $j^{*}_{U_{i}}(h^{k-1})$ in (1.3) and we get
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\!\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=r^{n-k+1}j^{*}_{U_{i}}(h^{n-k+1})\smile j^{*}_{U_{i}}(h^{k-1})\,$,
which implies
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=r^{n-k+1}j^{*}_{U_{i}}(h^{n})\,.$
For each $\alpha$ we have
$\widehat{W_{i\alpha}}]\smile h^{k-1}=P^{-1}[W_{i\alpha}]\smile P^{-1}[W_{i}]=P^{-1}\big([W_{i\alpha}\cap W_{i}]\big)=d_{i\alpha}\,P^{-1}[p_{i\alpha}],$
where $\{p_{i\alpha}\}=W_{i\alpha}\cap W_{i}$ is the (scheme-theoretic) intersection point of $W_{i\alpha}$ with a generic complete intersection
$W_{i}=H_{1}^{i}\cap\cdots\cap H_{n-k+1}^{i},\qquad H_{1}^{i},\dots,H_{n-k+1}^{i}\in H^{0}(X,\mathcal{O}_{X}(1)),$
and
$d_{i\alpha}=\operatorname{length}(W_{i\alpha}\cap W_{i})=(h^{k-1}\!\cdot W_{i\alpha})>0.$
Moreover, $p_{i\alpha}\in U_{i}$ can be defined by hyperplane sections through $p_{i\alpha}$:
$P^{-1}[p_{i\alpha}]=[\,\widetilde{H}^{(i,\alpha)}_{1}\cap\cdots\cap\widetilde{H}^{(i,\alpha)}_{n}\,]=h^{n},\qquad\widetilde{H}^{(i,\alpha)}_{\ell}\in H^{0}(X,\mathcal{O}_{X}(1))\ \text{generic}.$
Thus
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big(d_{i\alpha}P^{-1}[p_{i\alpha}]\big)=r^{n-k+1}\,j^{*}_{U_{i}}(h^{n}),$
and $j^{*}_{U_{i}}(h^{n})=j^{*}_{U_{i}}\big(P^{-1}[p_{i\alpha}]\big)\neq 0$ because $p_{i\alpha}\in Z_{i}\subset U_{i}$. Hence,
(1.4) $\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,d_{i\alpha}=r^{\,n-k+1}.$
Baum–Bott in cohomology gives
$c_{1}(N_{\mathscr{F}})^{\,n-k+1}=\sum_{i=1}^{s}\widehat{\mathrm{Res}}(\mathscr{F},Z_{i})=\sum_{i=1}^{s}\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,[\widehat{W_{i\alpha}}$].
Cupping with $h^{k-1}$ and using $\widehat{W_{i\alpha}}]\smile h^{k-1}=d_{i\alpha}\,h^{n}$ and (1.4), we obtain
$r^{\,n-k+1}h^{n}=c_{1}(N_{\mathscr{F}})^{\,n-k+1}\smile h^{k-1}=\sum_{i=1}^{s}\left(\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}d_{i\alpha}\,h^{n}\right)=\sum_{i=1}^{s}r^{\,n-k+1}h^{n}.$
Hence $h^{n}=s\,h^{n}$, and since $h$ is ample, $h^{n}\neq 0$; therefore $s=1$ and $Z$ is connected.
2. Proof Corollary 1
Theorem 1 guarantees that the curve $C$ is connected. Applying [6, Theorem 5], we conclude that
$h^{2}(\mathscr{F}\otimes\det(N_{\mathscr{F}})^{*})=c_{3}(\mathscr{F})\cap[X],$
and by [14, Proposition 3.6]
$c_{3}(\mathscr{F})\cap[X]=h^{0}(X,\mathscr{E}xt^{1}(\mathscr{F},\mathcal{O}_{X}))$
which is the number of isolated singularities of $\mathscr{F}$ counted with multiplicities, since by [5, Lemma 2.1], the foliation $\mathscr{F}$ has tangent sheaf locally free along $C$. Note that since $C$ is reduced, then it has no embedded points.
Now, suppose $X=\mathbb{P}^{3}$. Then, the dual of the determinant of the normal sheaf $N_{\mathscr{F}}$ is given by $\mathcal{O}_{\mathbb{P}^{3}}(-2-d),$ where $d$ denotes the degree of the foliation $\mathscr{F}$, see Remark 4.6. Applying Bott’s formulae [1, pg.12], we deduce
$h^{1}(\mathcal{O}_{\mathbb{P}^{3}})=h^{1}(T\mathbb{P}^{2}(-2-d))\simeq h^{1}(\Omega_{\mathbb{P}^{2}}^{1}(2-d))=0\quad\text{for all }d\geq 0.$
Furthermore, following the proof of Proposition 14 in [6], we obtain
$h^{2}(T\mathbb{P}^{3}\otimes\det(N_{\mathscr{F}})^{*})=h^{2}(T\mathbb{P}^{3}(-2-d))=h^{1}(\mathscr{I}_{C})=h^{0}(\mathcal{O}_{C})-1,$
which vanishes whenever $C$ is a reduced and connected curve.
3. Proof Corollary 2
According to [5, Theorem 3.11], a codimension one distribution $\mathscr{F}$ of positive degree on $\mathbb{P}^{3}$ with a locally free tangent sheaf and a smooth, connected singular scheme cannot be integrable. Remark 4.6 below states that Theorem 1 applies to this situation. Therefore, if $\mathscr{F}$ is integrable, then either $\mathscr{F}$ has degree zero and is a pencil of planes in which case ${\rm Sing}(\mathscr{F})$ is a line, or ${\rm Sing}(\mathscr{F})$ cannot be smooth.
4. Examples and remarks
In this section, we will present examples and observations that demonstrate the necessity of the hypotheses, in particular that Theorem 1 is sharp.
Example 4.1.
Let $X$ be the classical Hopf manifold of dimension $n$. This is obtained from $\mathbb{C}^{n}\setminus\{0\}$ dividing it by the group of automorphisms generated by a contraction, say $z\mapsto z/2$. It is clear that $X$ fibers over $\mathbb{P}^{n-1}$ with fiber an elliptic curve, the quotient of $\mathbb{C}^{*}$ by the contraction. So one has an elliptic fibration $X\mathrel{\mathop{\kern 0.0pt\to}\limits^{{\pi}}}\mathbb{P}^{n-1}$.
Consider now a 1-dimensional foliation $\mathscr{F}$ on $\mathbb{P}^{n-1}$ with isolated singularities. Its pull back $\pi^{\ast}\mathscr{F}$ is a 2-dimensional foliation on $X$ with singular set, a union of disjoint elliptic curves:
$\operatorname{Sing}(\pi^{*}\mathscr{F})=\bigsqcup_{p\in\operatorname{Sing}(\mathscr{F})}\pi^{-1}(p).$
If $\operatorname{Sing}(\mathscr{F})$ is reduced, then $\operatorname{Sing}(\pi^{*}\mathscr{F})$ has $d^{n-1}+\cdots+d+1$ connected components, see for instance [18, Theorem A].
In this example, neither the manifold in question is projective, nor is $\det(N_{\mathscr{F}})$ ample.
Example 4.2.
(Foliations on products) Let $\mathscr{F}$ be a holomorphic foliation of dimension $1$ with isolated singularities on a projective manifold $X$ of dimension $n$. If $Y$ is a projective manifold of dimension k1k-1, then we have a foliation $\pi_{2}^{*}\mathscr{F}$ on $X\times Y$ of dimension $k$, with singular set the disjoint union
$\operatorname{Sing}(\pi_{2}^{*}\mathscr{F})=\bigsqcup_{p\in\operatorname{Sing}(\mathscr{F})}\pi_{2}^{-1}(p),$
where $\pi_{2}^{-1}(p)=Y$ for all $p\in\operatorname{Sing}(\mathscr{F})$. Then $c_{1}(\det(N(\pi_{2}^{*}\mathscr{F})))=\pi_{2}^{*}c_{1}(\det(N_{\mathscr{F}}))$, and
$c_{1}(\det(N_{\mathscr{F}}(\pi_{2}^{*}\mathscr{F})))^{n+k-1}=0$
since n+k1>nn+k-1>n. So the determinant of the normal bundle is not ample.
The next example is a 2-dimensional foliation on $\mathbb{P}^{n}$, induced by the action of the affine group, whose singular set is the connected union of $n$ rational curves intersecting at a single point.
Example 4.3.
The next example generalizes to high dimensions an example from [4] for $n=3$; the foliation $\widetilde{\mathscr{F}}_{3,d}$ described below defines a rigid irreducible component of the space of foliations of codimension one on $\mathbb{P}^{3}$ of degree $d+1$, and $\widetilde{\mathscr{F}}_{3,1}$ generates the exceptional component of foliations of degree 2 [4, Cor. 3, p 1007].
In the affine space $\{[1:z]\}\simeq\mathbb{C}^{n}\subset\mathbb{P}^{n},\,n\geq 3$, let us denote $c_{k}(d)=c_{k}(T\mathbb{P}^{k}(d-1))$, and we define the complete vector fields
${u_{n,d}(z)=z_{1}\partial_{z_{1}}+c_{1}(d)\cdot z_{2}\partial_{z_{2}}+\cdots+c_{n-1}(d)\cdot z_{n}\partial_{z_{n}}}$${v_{n,d}(z)=\partial_{z_{1}}+c_{1}(d)\cdot z_{1}^{d}\partial_{z_{2}}+\cdots+c_{n-1}(d)\cdot z_{n-1}^{d}\partial_{z_{n}}.}$
Since $v_{n,d},u_{n,d}]=v_{n,d},$ they are the infinitesimal generators of an action of the affine group, $\mathfrak{aff}(\mathbb{C})\times\mathbb{C}^{n}\longrightarrow\mathbb{C}^{n}.$
The vector field $u_{n,d}$ is linear, it extends as a holomorphic section $\widetilde{u}_{n,d}$ of $T\mathbb{P}^{n}$. On the other hand, $v_{n,d}$ is polynomial of degree $d$, it extends as a holomorphic bundle map $\widetilde{v}_{n,d}:{\mathcal{O}}_{\mathbb{P}^{n}}(1-d)\longrightarrow T\mathbb{P}^{n}$, with a unique singularity at the point $p=[0:\cdots:0:1]\in\mathbb{P}^{n}$, with Milnor number
$\mu(\widetilde{v}_{n,d}(p))=c_{n}(T\mathbb{P}^{n}(d-1))=c_{n}(d).$
We define the foliation of dimension 2:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widetilde{\mathscr{F}}_{n,d}={\mathcal{O}}_{\mathbb{P}^{n}}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(1-d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{u}_{n,d}\oplus\widetilde{v}_{n,d}}$${T\mathbb{P}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\widetilde{\mathscr{F}}_{n,d}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
The singular set ${\rm Sing}(\widetilde{\mathscr{F}}_{n,d})=\Lambda_{1,d}\cup\Lambda_{2,d}\cup\dots\cup\Lambda_{n,d}$ is a union of rational curves $\Lambda_{j,d}$ of degrees $\deg(\Lambda_{1,d})=1$ and for $j>1,\quad\deg(\Lambda_{j,d})=c_{j-1}(d).$ Moreover
$\bigcap_{j=1}^{n}\Lambda_{j,d}=\{[0:\cdots:0:1]\},$
hence, ${\rm Sing}(\widetilde{\mathscr{F}}_{n,d})$ is connected. Notice that the restriction $\widetilde{\mathscr{F}}_{n,d}|_{\{Z_{n}=0\}}$ is $\widetilde{\mathscr{F}}_{n-1,d}$, and
${\rm Sing}(\widetilde{\mathscr{F}}_{n,d})={\rm Sing}(\widetilde{\mathscr{F}}_{n-1,d})\cup\Lambda_{n,d}.$
Denote $C_{k}=c_{k}(d)$, so the curve $\Lambda_{n,d}$ is parameterized by
$\Lambda_{n,d}([s:t])=[s^{C_{n-1}}:s^{(C_{n-1}-1)}t:s^{(C_{n-1}-C_{1})}t^{C_{1}}:\cdots:t^{C_{n-1}}].$
Remark 4.4.
Let $\mathscr{F}$ be a holomorphic foliation on $\mathbb{P}^{n}$ of dimension $k>1$, with $Z\subset{\rm Sing}(\mathscr{F})$ the non-empty component of dimension k1k-1. We can write
$Z=\bigcup_{j=1}^{s}Z_{j},$
where $Z_{j}$ are the connected components. In this case, we have
$Z_{i}\cap Z_{k}]=(d_{j}h^{n-k+1})\smile(d_{k}h^{n-k+1})=d_{j}d_{k}h^{2(n-k+1)}\in H^{2(n-k+1)}(\mathbb{P}^{n}),$
with $d_{j}=\deg(Z_{j})$ and $d_{k}=\deg(Z_{k})$. Thus, $Z_{i}\cap Z_{k}]\neq 0$ whenever $n\leq 2k-2$, implying that $Z$ is connected in this case. Consequently, the statement of Theorem 1 is not immediate in the case where n>2k2n>2k-2. Recently, in [10], it was shown that $\mathrm{dim}({\rm Sing}(\mathscr{F}))\geq k-1$, whenever $n\geq 2k$, meaning that $Z_{-}$ is empty.
Remark 4.5.
We note that the hypothesis that $k>1$ is also necessary: a generic foliation by curves on a complex projective manifold has 0-dimensional singularities consisting of several distinct points. In addition, [8] contains various examples of foliations by curves on $\mathbb{P}^{3}$ whose singular set is the union of two or more skew lines.
Remark 4.6.
Consider a holomorphic foliation $\mathscr{F}\subset TX$ of dimension $k$. By taking the maximal exterior power of the dual morphism
$\Omega^{1}_{X}\longrightarrow\mathscr{F}^{*},$
one obtains a morphism $\wedge^{k}\phi^{\vee}:\Omega^{k}_{X}\longrightarrow\det(\mathscr{F})^{*}.$ This, in turn, yields an induced morphism
$\Omega^{k}_{X}\otimes\det(\mathscr{F})\longrightarrow\mathcal{O}_{X},$
whose image is an ideal sheaf $\mathcal{I}_{Z}$ defining the singular set of the foliation $Z={\rm Sing}(\mathscr{F})\subset X$. Moreover, this construction determines a global section $\nu_{\mathscr{F}}\in H^{0}\left(X,\wedge^{k}TX\otimes\det(\mathscr{F})^{*}\right).$ By the isomorphism $\wedge^{k}TX\simeq\Omega_{X}^{n-k}\otimes\det(TX)$ and the relation $\det(N_{\mathscr{F}})\simeq\det(TX)\otimes K\mathscr{F},$ we conclude that the isomorphism
$\wedge^{k}TX\otimes K\mathscr{F}\simeq\Omega_{X}^{n-k}\otimes\det(TX)\otimes K\mathscr{F}\simeq\Omega_{X}^{n-k}\otimes\det(N_{\mathscr{F}})$
shows that the section $\nu_{\mathscr{F}}$ corresponds to a twisted (nk)(n-k)-form
$\omega_{\mathscr{F}}\in H^{0}\left(X,\Omega_{X}^{n-k}\otimes\det(N_{\mathscr{F}})\right)$
with values in $\det(N_{\mathscr{F}})$. A holomorphic foliation $\mathscr{F}$ of dimension $k$ on $\mathbb{P}^{n}$ can be described by a twisted (nk)(n-k)-form
$\omega\in H^{0}\left(\mathbb{P}^{n},\Omega_{\mathbb{P}^{n}}^{n-k}(d+n-k+1)\right),$
where $d\geq 0$ is referred to as the degree of the foliation $\mathscr{F}$. Then, $\det(N_{\mathscr{F}})=\mathcal{O}_{\mathbb{P}^{n}}(d+n-k+1)$ is ample. See for instance [9, Section 2.5].
Remark 4.7.
Observe that, a priori, each connected component is such that
$Z_{i}=\bigcup_{j=1}^{j(i)}Z_{ij}$
where $Z_{ij}$ are its irreducible components, for $i=1,\dots,s$. Theorem 1 implies the component of dimension k1k-1 of the singular set of the foliation $Z$ is connected. That is, $Z=Z_{i}$, for all $i$, and by Baum–Bott
$c_{1}(N_{\mathscr{F}})^{n-k+1}\cap[X]=\sum_{j=1}^{\ell}\lambda_{j}[Z_{j}]\in H_{2k-2}(Z),$
where $Z_{j}$ are the irreducible components of $Z$ and $\lambda_{j}$ are complex numbers that depend only on the local behavior of the leaves of $\mathscr{F}$ near $Z_{j}$.